OpenStudy (anonymous):

Which of the following is a polynomial with roots -3, -3i, and 3i ? x3 - 3x2 + 9x - 27 x3 - 9x2 + 9x - 81 x3 + 3x2 + 9x + 27 x3 + 9x2 + 9x + 81

OpenStudy (phi):

roots -3, -3i, and 3i means x= -3, x= -3i and x= 3i or x+3= 0 x+3i=0 or x-3i=0 or multiplying them out (x+3)(x+3i)(x-3i)= 0 can you multiply this out?

OpenStudy (anonymous):

what do you mean

OpenStudy (phi):

use FOIL to multiply (x+3i)(x-3i) then multiply that by (x+3)

OpenStudy (anonymous):

oh okay

OpenStudy (anonymous):

thanks